If f(x) is an even function in (-5,5) then the value of `X`not satisfying f(x)=f(x+1/x+2) is

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
f(x)=f(-x)
       Since f(x)= f((x+1)/(x+2))
Case1: x= (x+1)/(x+2)
            Solving we get x= (-1+√5)/2 , (-1-√5)/2
           Which lies within (-5,5)
Case 2: (x+1)/(x+2) = -x
x= (-3+√5)/2 , (-3-√5)/2
which lies within (-5,5)
Case 3: (x+1)/(x+2) = (-x+1)/(-x+2)
   Solving we get x=0
Which also lies within (-5,5)
Therefore no value of x that satisfies the given condition lies beyond (-5,5).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)